A magnetic field is perpendicular to the plane of a single-turn circular coil. The magnitude of the field is changing, so that an emf of 0.26 V and a current of 4.9 A are induced in the coil. The wire is then re-formed into a single-turn square coil, which is used in the same magnetic field (again perpendicular to the plane of the coil and with a magnitude changing at the same rate). What (a) emf and (b) current are induced in the square coil?
Answer #1

According to the question we need to find the emf and the current.

PART A) induced emf = dB/dt *A so dB/dt = Emf/A

Emf (square)/Emf (circle) = A square / A circle

Emf square = Emf circle * ( 3.14 / 4 ) = 0.26 *0.7854 = 0.2041V =====PART A)

PART B) I = V/R = 0.2041V / 0.053061224 = 3.8465A ============PART B)